a block of mass .50 kgs is moving with speed of 2 ms on smooth surface. it strikes with another mass of 1 kg then they move together as single body.  The energy losss during collision is 

Grade:11

1 Answers

Mayukh Chakrabarty
7 years ago
m=0.5 kg , u=2 m/sM=0.5+1=1.5 kgBy conservation of momentum,v=(mu)/M = 2/3Initial Kinetic Energy = (m * u^2)/2= 1Final Kinetic Energy = (M * v^2)/2 = 1/3So , loss = 2/3All data are in SI units .